You have 16-sided die and an 8-sided die. (Please provide an explanation/work for each response).

a) What are all the possible totals from rolling these 2 dice?

b) What's the probability that the total from rolling the 2 dice is 10?

c) What's the probability that the total from rolling the 2 dice is less than 7?
Answer #1

a) possible total from rolling these 2 dice are from 2 to 24 (total 23 values).

(as minimim sum can be 2 when both the dice shows 1 ; while maximum 24 will occur when first dice shows 16 and second dice shows 8)

b)total outcomes =16*8=128

number of outcome so that sum is 10 =8 {(2,8),(3,7),(4,6),(5,5),(6,4),(7,3),(8,2),(9,1)}

hence probability =8/128 =1/16

c)

number of ways sum is less then 7 =number of ways sum is 2+number of ways sum is 3+number of ways sum is 4+number of ways sum is 5+number of ways sum is 6 =1+2+3+4+5=15

hence probability =15/128